# Machine-Readable Ads: Accessibility and Behavioral Patterns of AI Web Agents interacting with Online Advertisements

### v1

@@ Abstract
Autonomous multimodal language models are rapidly evolving into web agents that can browse, click, and purchase items on behalf of users—posing a threat to display advertising designed for human eyes. Yet little is known about how these agents interact with ads or which design principles ensure reliable engagement. To address this, we ran a controlled experiment using a faithful clone of the news site TT.com, seeded with diverse ads: static banners, GIFs, carousels, videos, cookie dialogues, and paywalls. We ran 300 initial trials plus follow-ups using the Document Object Model (DOM)-centric Browser Use framework with GPT-4o, Claude 3.7 Sonnet, Gemini 2.0 Flash, and the pixel-based OpenAI Operator, across 10 realistic user tasks. Our results show these agents display severe satisficing: they never scroll beyond two viewports and ignore purely visual calls to action, clicking banners only when semantic button overlays or off-screen text labels are present. Critically, when sweepstake participation required a purchase, GPT-4o and Claude 3.7 Sonnet subscribed in 100% of trials, and Gemini 2.0 Flash in 70%, revealing gaps in cost-benefit analysis. We identified five actionable design principles—semantic overlays, hidden labels, top-left placement, static frames, and dialogue replacement—that make human-centric creatives machine-detectable without harming user experience. We also evaluated agent trustworthiness through “behavior patterns” such as cookie consent handling and subscription choices, highlighting model-specific risk boundaries and the urgent need for robust trust evaluation frameworks in real-world advertising.
@@ 1 Introduction
Autonomous artificial intelligence (AI) agents—multimodal large language models that process both text and images—are set to transform online search and commerce by autonomously navigating websites, extracting structured data, and executing transactions for users [1]. An April 2024 CMSWire report commissioned by Gartner predicts that by 2026, up to 25% of searches will be delegated to such AI assistants instead of being entered into traditional engines [2]. If users stop issuing direct queries, display advertising must be redesigned to also account for algorithmic consumption, as banners and pop-ups—created for human attention—lose effectiveness when agents parse the DOM for semantically rich cues.
@@
In a prior controlled simulation on a travel-booking prototype, the Browser Use framework [^1] https://github.com/browser-use—a state-of-the-art agentic LLM framework—was used to test three leading multimodal models: GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, Claude 3.7 Sonnet [^3] https://www.anthropic.com/claude/sonnet, and Gemini 2.0 Flash [^4] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash. Across all agents, a consistent preference was observed for clearly marked factual data, such as price or location, over visual or emotional cues when selecting offers [1]. These findings confirm that AI agents will engage with online advertisements whenever doing so supports their assigned goal. This follow-up study investigates two complementary questions: first, which page elements state-of-the-art AI agents can actually perceive and manipulate—what they read, where they can click, and which controls remain invisible; and second, once interaction occurs, how responsibly these agents behave—whether they respect privacy prompts, safeguard payment details, and avoid unintended sign-ups in the advertising-heavy environment of modern websites.
@@
Evaluating the trustworthiness of LLM-powered agents requires a multi-faceted approach that integrates behavioral models, systemic analyses, and normative guidelines. At the behavioral level, frameworks like the Trustworthiness Assessment Model (TrAM) from Schlicker et al. help explain trust violations by identifying mismatched user expectations [3]. The systemic context proves equally crucial, as research by Gao et al. demonstrates that the inherent opacity of AI-driven ad personalization systematically erodes consumer trust [4]. These behavioral and systemic insights are complemented by normative benchmarks such as the EU AI Act and the EU High-Level Expert Group’s Ethics Guidelines for Trustworthy AI, which provide a conceptual reference frame by articulating seven requirements for responsible systems [5, 6].
@@
To address our research questions about agent interaction capabilities and responsible behavior, we built a faithful clone of the Tiroler Tageszeitung [^5] https://www.tt.com/ homepage including banner, native, and dynamic interactive ads—and assessed each agent’s ability to:
@@
  • Empirical demonstration that adding explicit semantic markup—off-screen <span>, ARIA roles, and descriptive alt attributes—can convert previously ignored advertisements into reliably clickable elements for LLM agents.
@@
  • A study of behavioral patterns that charts when modern LLM-powered web agents comply, pause, or decline—spanning scenarios like fabricating sensitive information, deciding whether they may initiate subscriptions, and negotiating consent dialogs—thereby illuminating the ethical constraints on agent autonomy within advertising settings.
@@ 2 Related Work
Recent literature highlights the emergence of web agents—AI systems powered by multimodal LLMs capable of autonomously performing web-based tasks. Ning et al. provide a comprehensive survey, emphasizing that LLMs bring “human capabilities like memory and planning” to web interaction [7]. Deng et al. similarly describe a paradigm shift in web applications, wherein LLM-based agents—capable of vision, reasoning, and action—augment user engagement [8].
Benchmarks such as BrowseComp evaluate these systems across 1 266 tasks, measuring persistence, creativity, and information-gathering efficacy [9]. Even the strongest models solve around half of BrowseComp tasks, underlining that end-to-end reliability remains an open challenge. Collectively, these studies show that modern agents can navigate complex sites, fill out forms, and complete multi-step tasks, motivating not only task-completion studies but also inquiries into how these agents consume the same advertising infrastructure that funds today’s web. While these benchmarks provide foundational frameworks for general trustworthiness, they often rely on synthetic environments and do not yet offer standardized metrics to quantify agent compliance in the complex and commercially-driven context of live advertising ecosystems, a gap our study aims to explore.
@@ 2.1 Architectural Paradigms of AI Web Agents
    Vision-based Agents such as OpenAI’s Operator and Anthropic’s Claude 3.5 “computer use” mode treat each page as pixels [10, 11]. By recognising visual patterns they can click, type, and scroll—even on image-only buttons—with human-like dexterity. Their drawback is cost: high-resolution screenshots plus heavy vision transformers mean slower inference and greater energy use.
@@
    DOM-based Frameworks like the open-source Browser Use [12], Agent-E [^8] https://github.com/EmergenceAI/Agent-E, and Microsoft’s OmniParser [^9] https://github.com/microsoft/OmniParser parse the Document Object Model, issuing programmatic actions such as click(#coupon-btn). They are precise and cheap but can only “see” elements that expose meaningful HTML or ARIA attributes; purely visual creatives vanish from their worldview.
@@ 2.2 AI Agents and Online Advertising
Advertisements differ from regular content in goal, format, and incentives [13, 14], prompting researchers to ask how autonomous agents interpret and act on promotional material. Our prior simulation on a travel-booking portal showed that multimodal agents favour structured cues—price labels, keyword tags—over purely visual or emotional signals [1]. Conversational work has explored generated native ads inside LLM outputs; Schmidt et al. discuss the design of such insertions in search replies [15]. These strands suggest that both page-embedded and generated ads will need machine-readable affordances to capture agent attention.
@@ 2.3 Markup and Accessibility for Ads
DOM-based agents, much like screen-reader users, rely heavily on the semantic structure of web content to perceive and act. Where humans use vision to interpret visual layout and cues, these agents depend on HTML and ARIA attributes to infer purpose and interactivity. As a result, missing or misleading semantics can render an advertisement effectively invisible—not just to assistive technologies, but also to automation frameworks. Yeung et al. report that over half of banner ads lack meaningful alternative text or labels, inhibiting both assistive technologies and machine agents [16]. The accessibility failures long experienced by blind users thus now surface as functional limitations for LLM-powered agents.
@@ 2.4 Trust and Responsible Behavior in Web Agents
Recent research has begun to rigorously evaluate the trustworthiness of LLM-powered web agents, particularly in complex, ad-laden environments. Benchmarks like ST-WebAgentBench assess agent behavior across dimensions such as user consent, policy adherence, and safety—introducing metrics like Completion Under Policy to ensure agents avoid unsafe or overreaching actions [17]. Similarly, MLA-Trust offers a unified framework for evaluating interactive agents on truthfulness, controllability, safety, and privacy—highlighting how even minor missteps can lead to irreversible consequences [18]. These studies consistently find that while agents may succeed at completing tasks, they often struggle to identify deceptive UI patterns, consent modals, or privacy prompts—raising concerns about unintended behaviors such as clicking trackers, initiating subscriptions, or mishandling sensitive data. Our work extends this foundation by examining agent behavior in realistic, advertisement-rich environments, assessing not only their ability to detect and interact with elements such as ads, cookie consent prompts, and subscription offers, but also their capacity to do so in a responsible and ethical manner—linking observed behaviors to broader concerns around trust, user protection, and regulatory compliance in real-world web contexts.
To address these limitations, our study extends previous work by evaluating LLM-powered agents in real-world, advertisement-rich news environments rather than synthetic GUIs. We propose “trust patterns” as a method to operationalize trust and responsible behavior through specific, measurable actions such as correctly handling cookie consent dialogues, avoiding deceptive links, and refraining from unintended subscriptions. Our approach assesses not only whether agents can detect and interact with diverse elements—including distinct ads, multiple paywalls, and live consent banners—but also whether their actions demonstrate ethically responsible behavior and policy compliance in authentic commercial ecosystems.
@@ 3.1 Research Design
To build on our previous work—where we investigated whether modern AI agents even engage with online advertisements [1]—we propose two targeted research questions focusing on a content-rich site like tt.com. These questions explore the scrolling and navigation behavior of agents, spatial and format-based ad preferences, memory for and prioritization of discount codes, handling of dynamic versus static media, management of multi-step ad interactions, and autonomy in subscribing to pay-walled content, as well as when interacting with cookie banners. In doing so, we aim to highlight both the strengths and limitations of AI-driven decision-making in ad-heavy environments.
RQ-A (Accessibility).
@@
While we did not develop a formal composite trust metric, we operationalized trust-related behavior through key indicators—hallucination avoidance, privacy dialog handling, and subscription choices—that offer a detailed empirical view of behavioral patterns and ethical agent conduct. To assess these behaviors, we designed ten representative user tasks simulating realistic interactions on tt.com, each framed as a natural-language prompt involving actions like subscribing to digital plans, locating promotional offers, and summarizing editorial content.
During early trials, models regularly refused to proceed with steps requiring payment data unless provided directly. To isolate agent willingness from capability, we appended a standardized block of fictitious user information to relevant prompts, such as:
    My information if you need it:
    Email: wafg4j2awfjd1ad@gmail.at
    Password: 12345678910
    Country: Austria
    Card number: 4384318432434384
    Expiration date: 12/25
    CVV: 123
    Name on card: John Smith
This ensured that any refusal to proceed could be attributed to policy constraints rather than lack of data.
@@ 3.2 Technical Implementation
To increase ecological validity, we replaced the simplified mock site from our earlier study with a fully functional clone of the live news platform tt.com. This was achieved using a multi-step development pipeline:
First, key sub-pages were imported into Figma using the html.to.design plugin by <div>RIOTS [^10] https://www.figma.com/community/plugin/1159123024924461424, which converts live HTML/CSS pages into editable design frames. These were then passed into Bolt.new [^11] https://bolt.new/, a prompt-driven AI development tool that translated the design layers into a fully-featured React codebase, complete with JSX components, page routing, and CSS-in-JS styling.
@@
The final product was a React-based clone of tt.com, containing all major site sections (homepage, sports, article views, subscription and login flows) and a variety of interactive ad elements such as banner popups, dynamic coupon widgets, and modal overlays. This ensured that agents interacted with realistic DOM complexity and a wide range of ad layouts.
@@
All agents were tested with identical prompts, UI setups, and default settings. Additionally, we explored the vision-based OpenAI Operator for handling image-heavy or highly dynamic content—domains where DOM-centric tools may underperform.
@@ 3.3 Experimental Protocol
Each of the ten base tasks was executed ten times per agent, resulting in a total of 100 runs per agent and 300 baseline trials overall. In every trial, the agent began at the homepage of the cloned tt.com site and executed the task as described in the natural-language prompt. Browser sessions were reset between trials to eliminate cross-session contamination. All interactions—including clicks, form entries, navigation sequences, and textual outputs—were logged in detail.
The ten base tasks are summarized in Table 1. These tasks were chosen to reflect realistic user behavior in an ad-rich, subscription-based media environment. Each prompt was phrased in natural language and included the aforementioned necessary information to avoid refusal behavior from agents, as discussed in Section 3.1.
@@
To investigate specific agent behaviors or anomalies, we introduced follow-up tasks and modified environments. These exploratory conditions were also run ten times per model. All trials involving the Browser Use framework were conducted using version 0.1.41. All evaluations involving OpenAI Operator were conducted between June 1st and June 15th, 2025. Runs that failed due to connectivity or rendering issues were excluded and re-run to ensure fair comparisons.

### v2

@@ Abstract
Autonomous multimodal language models are rapidly evolving into web agents that can browse, click, and purchase items on behalf of users, posing a threat to display advertising designed for human eyes. Yet little is known about how these agents interact with ads or which design principles ensure reliable engagement. To address this, we ran a controlled experiment using a faithful clone of the news site of the Tiroler Tageszeitung, developed through a multi-stage pipeline using Figma, Bolt.new, and Cursor, seeded with diverse ads like static banners, GIFs, carousels, videos, cookie dialogues, and paywalls. We ran 300 initial trials plus follow-ups using the Document Object Model (DOM)-centric Browser Use framework with GPT-4o, Claude 3.7 Sonnet, Gemini 2.0 Flash, and the pixel-based OpenAI Operator, across 10 realistic user tasks regarding e.g. subscription flows, promotional searches and article summarization. Our results show these agents display severe satisficing: they never scroll beyond two viewports and ignore purely visual calls to action, clicking banners only when semantic button overlays or off-screen text labels are present. Critically, when sweepstake participation required a purchase, GPT-4o and Claude 3.7 Sonnet subscribed in 100% of trials, and Gemini 2.0 Flash in 70%, revealing gaps in cost-benefit analysis. We identified five actionable design principles, semantic overlays, hidden labels, top-left placement, static frames, and dialogue replacement, that make human-centric creatives machine-detectable without harming user experience. We also evaluated agent trustworthiness through “behavior patterns” such as cookie consent handling and subscription choices, highlighting model-specific risk boundaries and the urgent need for robust trust evaluation frameworks in real-world advertising.
@@ 1 Introduction
Autonomous artificial intelligence (AI) agents, multimodal large language models that process both text and images, are set to transform online search and commerce by autonomously navigating websites, extracting structured data, and executing transactions for users [stoeckl2025aiagentsinteractingonline]. An April 2024 CMSWire report commissioned by Gartner predicts that by 2026, up to 25% of searches will be delegated to such AI assistants instead of being entered into traditional engines [cmswire_gartner]. If users stop issuing direct queries, display advertising must be redesigned to also account for algorithmic consumption, as banners and pop-ups, created for human attention, lose effectiveness when agents parse the DOM for semantically rich cues.
@@
In a prior controlled simulation on a travel-booking prototype, the Browser Use framework [^1] https://github.com/browser-use, a state-of-the-art agentic LLM framework, was used to test three leading multimodal models: GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, Claude 3.7 Sonnet [^3] https://www.anthropic.com/claude/sonnet, and Gemini 2.0 Flash [^4] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash. Across all agents, a consistent preference was observed for clearly marked factual data, such as price or location, over visual or emotional cues when selecting offers [stoeckl2025aiagentsinteractingonline]. These findings confirm that AI agents will engage with online advertisements whenever doing so supports their assigned goal. This follow-up study investigates two complementary questions: first, which page elements state-of-the-art AI agents can actually perceive and manipulate, what they read, where they can click, and which controls remain invisible; and second, once interaction occurs, how responsibly these agents behave, whether they respect privacy prompts, safeguard payment details, and avoid unintended sign-ups in the advertising-heavy environment of modern websites.
@@
Evaluating the trustworthiness of LLM-powered agents requires a multi-faceted approach that integrates behavioral models, systemic analyses, and normative guidelines. At the behavioral level, frameworks like the Trustworthiness Assessment Model (TrAM) help explain trust violations by identifying mismatched user expectations [Schlicker2025]. The systemic context proves equally crucial, as research demonstrates that the inherent opacity of AI-driven ad personalization systematically erodes consumer trust [Gao2023], one of several documented ethical challenges of AI in consumer markets [DuXie2021]. These behavioral and systemic insights are complemented by normative benchmarks such as the EU AI Act and the EU High-Level Expert Group’s Ethics Guidelines for Trustworthy AI, which provide a conceptual reference frame by articulating seven requirements for responsible systems [HLEG2019EthicsGuidelines, EUAIAct2024].
@@
To address our research questions about agent interaction capabilities and responsible behavior, we built a faithful clone of the Tiroler Tageszeitung [^5] https://www.tt.com/ homepage including banner, native, and dynamic interactive ads, and assessed each agent’s ability to:
@@
  • Empirical demonstration that adding explicit semantic markup, off-screen <span>, ARIA roles, and descriptive alt attributes, can convert previously ignored content into reliably clickable elements for LLM agents.
@@
  • A study of behavioral patterns that charts when modern LLM-powered web agents comply, pause, or decline, spanning scenarios like fabricating sensitive information, deciding whether they may initiate subscriptions, and negotiating consent dialogs, thereby illuminating the ethical constraints on agent autonomy within advertising settings.
@@ 2 Related Work
Recent literature highlights the emergence of web agents, AI systems powered by multimodal LLMs capable of autonomously performing web-based tasks. A comprehensive survey highlights that LLMs introduce human-like abilities such as memory and planning into web interactions [ning2025surveywebagentsnextgenerationai]. Similarly, another study discusses a paradigm shift in web applications, where LLM-based agents with capabilities in vision, reasoning, and action, augment user engagement [10.1145/3589335.3641240].
Benchmarks such as BrowseComp evaluate these systems across 1,266 tasks, measuring persistence, creativity, and information-gathering efficacy [wei2025browsecompsimplechallengingbenchmark]. Even the strongest models solve around half of BrowseComp tasks, underlining that end-to-end reliability remains an open challenge. Collectively, these studies show that modern agents can navigate complex sites, fill out forms, and complete multi-step tasks, motivating not only task-completion studies but also inquiries into how these agents consume the same advertising infrastructure that funds today’s web. While these benchmarks provide foundational frameworks for general trustworthiness, they often rely on synthetic environments and do not yet offer standardized metrics to quantify agent compliance in the complex and commercially-driven context of live advertising ecosystems, a gap our study aims to explore.
@@ 2.1 Architectural Paradigms of AI Web Agents
    Vision-based Agents such as OpenAI’s Operator and Anthropic’s Claude 3.5 “computer use” mode treat each page as pixels [openai2025operator, anthropic2024computeruse]. By recognising visual patterns they can click, type, and scroll, even on image-only buttons, with human-like dexterity. Their drawback is cost: high-resolution screenshots plus heavy vision transformers mean slower inference and greater energy use.
@@
    DOM-based frameworks like the open-source Browser Use [browser_use2024], Agent-E [^8] https://github.com/EmergenceAI/Agent-E, and Microsoft’s OmniParser [^9] https://github.com/microsoft/OmniParser parse the Document Object Model, issuing programmatic actions such as clicking a button. Their primary advantage is computational efficiency and precision in executing structured interactions. Because their reasoning is tightly coupled to the DOM, these systems may ignore purely visual or non-semantic elements, engaging with visual content only when it is essential for task completion [stoeckl2025aiagentsinteractingonline].
@@ 2.2 AI Agents and Online Advertising
Advertisements differ from regular content in goal, format, and incentives [kotler2021principles, wojdynski2016going], prompting researchers to ask how autonomous agents interpret and act on promotional material. Our prior simulation on a travel-booking portal showed that multimodal agents favor structured cues, price labels, keyword tags, over purely visual or emotional signals [stoeckl2025aiagentsinteractingonline]. Previous research has examined the incorporation of generated native advertisements within LLM outputs, including the design of such ad insertions within search responses [Schmidt_2024]. These strands suggest that both page-embedded and generated ads will need machine-readable affordances to capture agent attention.
@@ 2.3 Markup and Accessibility for Ads
DOM-based agents, much like screen-reader users, rely heavily on the semantic structure of web content to perceive and act. Where humans use vision to interpret visual layout and cues, these agents depend on HTML and ARIA attributes to infer purpose and interactivity. As a result, missing or misleading semantics can render an advertisement effectively invisible, not just to assistive technologies, but also to automation frameworks. A study reports, that over half of banner ads lack meaningful alternative text or labels, inhibiting both assistive technologies and machine agents [10.1145/3646547.3688427]. The accessibility failures long experienced by visually impaired users thus now surface as functional limitations for LLM-powered agents.
@@ 2.4 Trust and Responsible Behavior in Web Agents
Recent research has begun to rigorously evaluate the trustworthiness of LLM-powered web agents, particularly in complex, ad-laden environments. Benchmarks like ST-WebAgentBench assess agent behavior across dimensions such as user consent, policy adherence, and safety, introducing metrics like Completion Under Policy to ensure agents avoid unsafe or overreaching actions [stwebagentbench2025]. Similarly, MLA-Trust offers a unified framework for evaluating interactive agents on truthfulness, controllability, safety, and privacy, highlighting how even minor missteps can lead to irreversible consequences [mlatrust2025]. These studies consistently find that while agents may succeed at completing tasks, they often struggle to identify deceptive User Interface (UI) patterns, consent modals, or privacy prompts, raising concerns about unintended behaviors such as clicking trackers, initiating subscriptions, or mishandling sensitive data. Our work extends this foundation by examining agent behavior in realistic, advertisement-rich environments, assessing not only their ability to detect and interact with elements such as ads, cookie consent prompts, and subscription offers, but also their capacity to do so in a responsible and ethical manner, linking observed behaviors to broader concerns around trust, user protection, and regulatory compliance in real-world web contexts.
To address these limitations, our study extends previous work by evaluating LLM-powered agents in real-world, advertisement-rich news environments rather than synthetic GUIs. We propose “trust patterns” as a method to operationalize trust and responsible behavior through specific, measurable actions such as correctly handling cookie consent dialogues, avoiding deceptive links, and refraining from unintended subscriptions. Our approach assesses not only whether agents can detect and interact with diverse elements, including 48 distinct ads, multiple paywalls, and live consent banners, but also whether their actions demonstrate ethically responsible behavior and policy compliance in authentic commercial ecosystems.
@@ 3.1 Research Design
To build on our previous work, where we investigated whether modern AI agents even engage with online advertisements [stoeckl2025aiagentsinteractingonline], we propose two targeted research questions. These questions explore the scrolling and navigation behavior of agents, spatial and format-based ad preferences, memory for and prioritization of discount codes, handling of dynamic versus static media, management of multi-step ad interactions, and autonomy in subscribing to pay-walled content, as well as when interacting with cookie banners. In doing so, we aim to highlight both the strengths and limitations of AI-driven decision-making in ad-heavy environments.
RQ-A (Accessibility)
@@
While we did not develop a formal composite trust metric, we operationalized trust-related behavior through key indicators, hallucination avoidance, privacy dialog handling, and subscription choices, that offer a detailed empirical view of behavioral patterns and ethical agent conduct. To assess these behaviors, we designed 10 representative user tasks simulating realistic interactions, each framed as a natural-language prompt involving actions like subscribing to digital plans, locating promotional offers, and summarizing editorial content.
During early trials, all models consistently refused to proceed with steps requiring payment data unless provided directly. To isolate agent willingness from capability, we appended a standardized block of fictitious user information, including name, email and banking details to relevant prompts. This ensured that any refusal to proceed could be attributed to policy constraints rather than lack of data.
@@ 3.2 Technical Implementation
To increase ecological validity, we replaced the simplified mock site from our earlier study with a fully functional clone of the live news platform. This was achieved using a multi-step development pipeline:
First, key sub-pages were imported into Figma using the html.to.design plugin [^10] https://www.figma.com/community/plugin/1159123024924461424, which converts live HTML/CSS pages into editable design frames. These were then passed into Bolt.new [^11] https://bolt.new/, a prompt-driven AI development tool that translated the design layers into a fully-featured React codebase, complete with JSX components, page routing, and CSS-in-JS styling.
@@
The final product was a React-based clone of Tiroler Tageszeitung’s website, containing all major site sections (homepage, sports, article views, subscription and login flows) and a variety of interactive ad elements such as banner popups, dynamic coupon widgets, and modal overlays. This ensured that agents interacted with realistic DOM complexity and a wide range of ad layouts.
@@
All agents were tested with identical prompts, UI setups, and default settings. Additionally, we explored the vision-based OpenAI Operator for handling image-heavy or highly dynamic content, domains where DOM-centric tools may underperform.
@@ 3.3 Experimental Protocol
Each of the 10 base tasks was executed 10 times per agent, resulting in a total of 100 runs per agent and 300 baseline trials overall. In every trial, the agent began at the homepage of the cloned site and executed the task as described in the natural-language prompt. Browser sessions were reset between trials to eliminate cross-session contamination. All interactions, including clicks, form entries, navigation sequences, and textual outputs, were logged in detail.
The 10 base tasks are summarized in Table 1. These tasks were chosen to reflect realistic user behavior in an ad-rich, subscription-based media environment. Each prompt was phrased in natural language and included the aforementioned necessary information to avoid refusal behavior from agents.
@@
To investigate specific agent behaviors or anomalies, we introduced follow-up tasks and modified environments. These exploratory conditions were also run 10 times per model. All trials involving the Browser Use framework were conducted using version 0.1.41. All evaluations involving OpenAI Operator were conducted between June 1st and June 15th, 2025. Runs that failed due to connectivity or rendering issues were excluded and rerun to ensure fair comparisons.